OpenStudy (anonymous):

The value of x when 2^(3x+3) = 2^(3x+1) + 48 is??

OpenStudy (ajprincess):

Assume 3x+1=y 2^(y+2)=2^y*2^2 =4*2^y 4*2^y=2^y+48 3*2^y=48 2^y=16 2^y=2^4 y=4 3x+1=4 3x=3 x=1

OpenStudy (ajprincess):

did u get t?

OpenStudy (anonymous):

no how 2^(y+2)=2^y*2^2???

OpenStudy (anonymous):

ur answer is correct

OpenStudy (ajprincess):

as u can see I assumed 3x+1 as y. 3x+3=3x+1+2 =y+2 Upto this is t clear?

OpenStudy (anonymous):

where that 2 came from

OpenStudy (ajprincess):

3x+3 can be written as 3x+1+2. do u get this?

OpenStudy (ajprincess):

@shameer1

OpenStudy (anonymous):

ok plzz go on

OpenStudy (ajprincess):